OpenStudy (anonymous):

Can someone help me with a homework problem? Simplifying complex fractions with variables. I want to learn HOW to do it, not just get the answer- I have a test tomorrow.

OpenStudy (anonymous):

I can try!

OpenStudy (anonymous):

\[\frac{ \frac{ x }{ x+2 } }{ \frac{ 1 }{ x }+\frac{ 1 }{ x+2}}\]

OpenStudy (anonymous):

I have no idea how to even begin :P

OpenStudy (anonymous):

It is kind hard to see :D

OpenStudy (anonymous):

*kinda

OpenStudy (anonymous):

The equation you mean?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):